A construction company in Naples, Florida, is struggling to sell condominiums. In order to attract buyers, the company has made numerous price reductions and better financing offers. Although condominiums were once listed for $350,000, the company believes that it will be able to get an average sale price of $255,000. Let the price of these condominiums in the next quarter be normally distributed with a standard deviation of $18,000. Use Table 1.

a.

What is the probability that the condominium will sell at a price (i) Below $244,000?, (ii) Above $277,000? (Round "z" value to 2 decimal places and final answer to 4 decimal places.)

   Probability
  Below $244,000      
  Above $277,000      
b.

The company is also trying to sell an artist’s condo. Potential buyers will find the unusual features of this condo either pleasing or objectionable. The manager expects the average sale price of this condo to be the same as others at $255,000, but with a higher standard deviation of $25,000. What is the probability that this condo will sell at a price (i) Below $244,000?, (ii) Above $277,000? (Round your answers to 4 decimal places.)

    Probability
  Below $244,000      
  Above $277,000      

rev: 08_01_2013_QC_33082
Answer #1

a) i) P(X < 244000)

= P(Z < -0.61)

= 0.2709

ii) P(X > 277000)

= P(Z > 1.22)

= 1 - P(Z < 1.22)

= 1 - 0.8888

= 0.1112

b) i) P(X < 244000)

= P(Z < -0.44)

= 0.3300

ii) P(X > 277000)

= P(Z > 0.88)

= 1 - P(Z < 0.88)

= 1 - 0.8106

= 0.1894